A is now three times older than B. In 5 years A will be twice as old as B will be then. Find their present ages.
I am an English teacher tutoring math and can't remember how to do this one.

Found 2 solutions by rothauserc, Alan3354:
Answer by rothauserc(4718)

A = 3B
A +5 = 2(B +5)
A +5 = 2B +10
now substitute for A = 3B
3B +5 = 2B +10
subtract 2B form both sides of =
B +5 = 10
B = 5 and A = 15
****************************
check answer by substituting for A and B in the second equation
15 +5 = 2(5+5)
20 = 20
our answer checks

Answer by Alan3354(69443)

I am an English teacher...
============
Then you probably know that 3 times older = 4 times as old. Right?